# NeuroIDBench: An Open-Source Benchmark Framework for the Standardization of Methodology in Brainwave-based Authentication Research

### v3

@@ 3 NeuroIDBench
As illustrated in Figure 2, NeuroIDBench is organized into five integral components: dataset, preprocessing, feature extraction, evaluation, and analysis. Our benchmarking workflow is influenced by MOABB (Mother of all BCI benchmarks) jayaram2018moabb work. In line with our research aims, we have tailored and refined their methodology to create a customized benchmarking suite specifically designed for evaluating and analyzing brainwave authentication systems. In the following, we detail NeuroIDBench’s components, justifying the design choices.
@@ 3.1 Dataset Loading
In choosing our dataset, we focused on Event-Related Potential (ERP) paradigms, particularly the P300 and N400, which are the most well-known and easier to elicit.The P300 reaction is triggered with the ‘oddball’ paradigm, i.e., including a rare stimulus in a sequence picton1992p300 , while the N400 is linked to semantic processing debruille1996n400 . Our selection methodology involved a thorough review of publicly accessible datasets utilizing these ERP paradigms. We carefully examined and assessed over 29 datasets (see Table 2). Ultimately, we chose 9 datasets for further investigation. This selection included two multi-session datasets to examine performance over multiple sessions cogbci ; lee2019eeg , five datasets with over 30 subjects to evaluate performance in datasets with a relatively large number of samples mantegna2019distinguishing ; erpcore ; brainInvaders15a ; won2022eeg , and two datasets with just 13 subjects to explore performance in smaller datasets hubner2017learning ; sosulski2019spatial .
@@ 3.2 Pre-processing
In this module, we conduct pre-processing on the standardized MNE Raw data. Various methods exist for cleansing artifacts; however, the procedures must remain consistent to ensure the validity of comparisons between algorithms or datasets jayaram2018moabb . We adhere to established best practices commonly employed in pre-processing methodologies within brainwave authentication studies gui2019survey . To cleanse EEG artifacts, we begin by implementing bandpass filtering in the range of 1 to 50 Hz. Subsequently, we extract samples from the raw signals, temporally aligning the data to a range spanning from -200 to 800 ms relative to the onset of the stimulus. Afterward, Following the baseline correction, where we subtract the mean of a baseline period (-200 to 0 ms) from the entire sample data points to mitigate drift effects like DC offsets fallahi2023brainnet . Finally, we employ a peak-to-peak rejection method to eliminate large artifacts attributed to eye or muscle movements from contaminating the EEG data.
@@ 3.3 Feature Extraction
To compute AR coefficients, our system fits AR models to the pre-processed samples, employing a one-second duration of time series data arias2023performance . The estimation of these AR coefficients utilized the Yule-Walker method pardey1996review . Determining the ideal order for AR modeling poses a challenge, as higher orders escalate computational demands, while exceedingly low orders inadequately represent the signal zhang2017classification . To address this, our system allows users the flexibility to define their preferred AR order, enabling a tailored approach aligned with their specific requirements.
@@
Furthermore, NeuroIDBench implements a deep learning based approach using a TNN with triplet loss function to transform time-series EEG data into compact brain embeddings. Building on the work of Fallahi et al. fallahi2023brainnet , our TNN structure integrates three Convolutional Neural Network (CNN) branches, each featuring five convolutional layers, as illustrated in Figure 3. These branches received 1-second samples as input, structured in a two-dimensional array where rows correspond to channel indices, and columns represent discrete-time measurements. After each convolutional layer, we applied an average pooling layer to reduce the input vectors’ dimensionality while preserving the unique characteristics of each brainwave.
Figure 3: The CNN’s network architecture within the Twin neural Network is designed to produce a condensed 32-bit embedding from brain data samples, serving as an efficient method to compute latent representations of inputs fallahi2023brainnet .
@@ 3.4 Evaluation
Within this module, we actively utilize both learning methods for authentication, namely supervised learning employing shallow classifiers and similarity-based learning through Twin Neural Networks. Furthermore, we engage in classification under two distinct evaluation strategies: single-session and multi-session evaluation:
@@
For the known attacker scenario, we implemented a Stratified cross-validation approach with k=4 to divide the single-session data into training and testing sets. This choice ensures the representation of features from both classes in both training and testing data across each fold. We take the step of eliminating users with fewer than four samples from the datasets to guarantee sufficient samples for both training and testing arias2023performance . Conversely, the unknown attacker approach adopts the GroupKFold cross-validation strategy with k=4, where grouping is based on SubjectID, resulting in non-overlapping training and testing sets for users in each cross-validation round. Evaluation metric results are aggregated across all folds, providing averaged and comprehensive reports. In datasets like COGBCI and Lee2019, which had multiple sessions, we evaluate each session separately, subsequently averaging the results from all the sessions as single-session evaluation results. Multi-session evaluation, described later in this section.
Additionally, we actively apply z-scaler normalization to both training and testing data to address potential issues arising from variations in feature scales, ensuring equitable contribution from all features during model training hastie2009elements . This approach aligns with established best practices in machine learning pedregosa2011scikit , reinforcing the reliability and generalization capability of our EEG-based authentication system.
Multi-Session Evaluation: The multi-session evaluation methodology was applied to two EEG datasets, namely COGBCI cogbci and Lee2019 lee2019eeg , each consisting of three and two sessions, respectively. To address potential overfitting, a cross-validation strategy was employed. This involved dividing the attacker subjects for the shallow classifiers and dividing the subject of latent space learning and evaluation.
@@ 3.5 Analysis
This module provides methodologies for conducting statistical analyses on the performance of various datasets and algorithms jayaram2018moabb . Following the evaluation process, the performance of authentication methods across each cross-validation fold is systematically analyzed by computing essential evaluation metrics, EER serves as a summary metric, indicating the point where the False Match Rate (FMR) and the False Non-Match Rate (FNMR) are equal. Additionally, we report FNMR at specific FMR thresholds of 1%, 0.1%, and 0.01%, allowing for a granular examination of the model’s performance under different operating conditions and security requirements. The FMR represents the likelihood of an unauthorized user being mistakenly authenticated by the system, reflecting the system’s vulnerability to a zero-effort attack. The objective is to achieve a low FMR while maintaining a reasonable FNMR. A high FNMR can result in legitimate users being repeatedly denied access, negatively impacting the device’s usability. Hence, it is crucial to strike an optimal balance between FMR and FNMR to ensure robust security without undermining user experience.
By adopting this approach, we align our evaluation strategy with the foundational principles outlined in Jain et al. jain1996introduction and Ross at al. ross2006handbook . This analytical rigor ensures that the performance evaluation is not only robust but also reflective of the diverse operational scenarios encountered in real-world applications.
@@ 4.1 Preprocessing
Figure 4 presents the EER results for the sample duration experiment. The results indicate that longer samples may enhance outcomes in shallow classifiers, but no significant improvement was observed in the case of Twin neural networks. It is important to note that for longer samples (e.g., 2 seconds), it was necessary to downsample the sample to mitigate the increased data dimensionality, a step taken due to hardware limitations as suggested in source code of BrainNet fallahi2023brainnet paper as well.
@@ 4.3 Known and UnKnown Attacker Approach
The known attacker model represents an unrealistic scenario; however, given its existence in some previous studies, we investigated the performance gap between known and unknown attacker scenarios. Figure 7 displays the EER for these two scenarios, based on an sample duration of 1 second to ensure comparability with previous work. This analysis was conducted without sample rejection and utilized PSD combined with AR order 1 as the feature extraction method for shallow classifiers.
@@
For the practical application of brainwave authentication, it is essential to develop a model capable of handling multi-session authentication. However, there is a notable limitation in the availability of datasets for multi-session studies. Our investigation focuses on comparing single-session and multi-session authentication to understand the performance gap between these two approaches. We expect an increase in EER for multi-session authentication, as it represents a more complex challenge. This increase in difficulty is attributed to additional noise factors inherent in multi-session settings. Additionally, variations in EEG electrode placement across sessions and possible changes in brain states add to the complexity, making multi-session authentication a more challenging task compared to single-session scenarios.
@@ 4.5 Shallow Classifiers vs Similarity-Based Approach
2.44
12.92
5.14
4.45
0.46
0.74
0.43
BrainInvaders15a
Single-Session
FNMR at 1% FMR
19.87
71.26
43.39
31.55
0.69
2.06
2.12
FNMR at 0.1% FMR
34.62
83.22
58.55
38.77
1.47
4.54
3.70
@@
5.72
4.75
3.41
10.16
0.90
2.73
0.81
ERPCORE_N400
Single-Session
FNMR at 1% FMR
22.70
35.83
25.67
76.06
1.96
7.54
1.17
@@
6.87
5.73
3.45
9.11
1.37
3.68
1.53
ERPCORE_P300
Single-Session
FNMR at 1% FMR
26.65
39.28
30.03
59.55
4.36
10.47
4.30
@@
3.28
1.58
0.84
5.89
0.36
0.55
0.03
Huebner_LLP
Single-Session
FNMR at 1% FMR
26.48
8.41
5.06
34.01
0.31
0.51
0.02
@@ 4.6 Comparison with Related Work
The comparison of our results will initially focus on the results published by Arias et al. arias2021inexpensive and Fallahi et al. fallahi2023brainnet . This is due to the establishment of our benchmark, which is based on their implementations. Moreover, both studies utilized the ERPCORE erpcore P300 and N400 datasets, similar to our approach. Arias et al. arias2021inexpensive investigate the performance of different shallow classifiers within single-session brainwave authentication. The results indicate 1.9% EER for N400 and 3% EER for P300 paradigm with a sample rejection of 120 µV and RF classifier, we observe the same results where we achieved 2.9% EER for P300 and 1.13% EER to N400 paradigm with the sample rejection of 100 µV and RF classifier. Also, Fallahi et al. fallahi2023brainnet acquired 2.01% and 1.37% respectively for P300 and N400 ERPCORE and 0.14 for P300:bi2015a dataset bi2015a , where we obtained %1.53, 0.83% , and 0.43 respectively. As can be observed the results are fairly similar and trends are the same. The small difference in results could be from a different random seed, version of libraries, and some evaluation parameters like epoch rejection rate for shallow classifiers and number of epochs for twin neural networks. To enhance the reproducibility of our result, we provided a docker container in the GitHub of NeuroIDBench.
@@
Preprocessing: Our results about the impact of sample duration corroborate prior research on shallow classifiers, which indicates that longer sample durations can result in reduced EER or increased accuracy arias2021inexpensive ; tran2019eeg ; carrion2019method ; suppiah2018biometric . However, the current version of the twin neural network requires modification to effectively utilize longer samples.
@@
Feature Extraction: The results of our feature extraction analysis demonstrate that PSD features notably outperform AR features, corroborating findings from Huang et al. huang2022m3cv . However, there is a lack of consensus in the literature regarding the optimal order for AR features. Studies by Arias et al. arias2023performance and Brigham brigham2010subject suggest that lower AR orders yield better results, whereas research by Kaewwit et al. kaewwit2017high and Zhang et al. zhang2017classification indicates improved outcomes with higher orders. Our findings align with the former, showing that lower-order AR features are more effective in 8 out of 9 datasets analyzed.
@@
Known and Unkown attacker: In exploring scenarios involving known and unknown attackers, Wu et al. wu2018eeg conducted research but found no notable performance decline in the unknown attacker scenario, and even reported an improvement. Similarly, Panzino et al. panzino2022eeg studied the same issue and observed only a 5% decrease in accuracy. Contrarily, Arias et al.arias2023performance and Fallahi et al.fallahi2023brainnet reported an increase in Equal Error Rate (EER) ranging from 10% to 17.5 times. Our results indicate that the unknown attacker scenario is consistently more effective than the known attacker scenario, although the extent of this difference varies across different datasets and models. An average calculation across nine datasets and seven models shows that the EER for unknown attackers is 60.2% higher than known attacker scenarios.
@@
Single and Multi-session: Multi-session analysis is crucial for the practical application of brainwave authentication, yet it presents heightened challenges compared to single-session scenarios. In multi-session environments, the availability of public datasets is notably scarcer, and the practice of open-source sharing becomes more constrained. Results from various studies, such as Maiorana maiorana2021learning , demonstrate reasonable EER as low as 4.8%. Seha et al. seha2019new also reported a marginal difference of less than 0.5% EER between single and multi-session scenarios under specific feature extraction methods and LDA classifiers. Furthermore, Wu et al. wu2018eeg observed a 1.3% improvement in false rejection rate performance in the second session compared to the first session in single-session scenarios. In contrast, our research reveals a significant disparity between single-session and multi-session results, with our best outcome showing an 11.99% EER. This finding aligns with Huang et al. huang2022m3cv , who also reported a substantial difference between within-session and cross-session evaluations. However, it is important to note that the studies by Maiorana maiorana2021learning , Seha et al. seha2019new , and Wu et al. wu2018eeg do not provide public datasets or open-source code for other researchers to utilize and build upon. Therefore, there is an urgent need for more open-source and publicly available datasets in multi-session studies in this field. In fact, understanding the exact state of this challenge is crucial to determine what steps are necessary to advance the field.
@@
Shallow Classifiers vs Similarity-Based Approach: The current trend in biometric authentication leans towards deep learning approaches. We observed the same trend in brainwave authentication where several papers based on TNN fallahi2023brainnet ; maiorana2021learning ; schons2018convolutional or CNN bidgoly2022towards to improve authentication. However, a comparison between shallow classifiers and deep learning methods is essential. This is because the scarcity of data can hinder the training of deep learning networks, particularly in managing the high variation in brainwave data. Several studies debie2021session ; fallahi2023brainnet demonstrate that deep learning approaches can outperform shallow classifiers. However, our large-scale benchmark emphasizes, despite the impressive results of TNN, they do not consistently surpass shallow classifies, Notably, in multi-session scenarios, TNNs underperform, potentially due to insufficient data to accommodate the high variability between sessions.

### v5

@@ 3 NeuroIDBench
As illustrated in Figure 2, NeuroIDBench is organized into five integral components: dataset, preprocessing, feature extraction, evaluation, and analysis. Our benchmarking workflow is influenced by MOABB (Mother of all BCI benchmarks) [26] work. In line with our research aims, we have tailored and refined their methodology to create a customized benchmarking suite specifically designed for evaluating and analyzing brainwave authentication systems. In the following, we detail NeuroIDBench’s components, justifying the design choices.
@@ 3.1 Dataset Loading
In choosing our dataset, we focused on Event-Related Potential (ERP) paradigms, particularly the P300 and N400, which are the most well-known and easier to elicit.The P300 reaction is triggered with the ‘oddball’ paradigm, i.e., including a rare stimulus in a sequence [67], while the N400 is linked to semantic processing [68]. Our selection methodology involved a thorough review of publicly accessible datasets utilizing these ERP paradigms. We carefully examined and assessed over 29 datasets (see Table 2). Ultimately, we chose 9 datasets for further investigation. This selection included two multi-session datasets to examine performance over multiple sessions [49, 61], five datasets with over 30 subjects to evaluate performance in datasets with a relatively large number of samples [43, 36, 56, 66], and two datasets with just 13 subjects to explore performance in smaller datasets [59, 60].
@@ 3.2 Pre-processing
In this module, we conduct pre-processing on the standardized MNE Raw data. Various methods exist for cleansing artifacts; however, the procedures must remain consistent to ensure the validity of comparisons between algorithms or datasets [26]. We adhere to established best practices commonly employed in pre-processing methodologies within brainwave authentication studies [4]. To cleanse EEG artifacts, we begin by implementing bandpass filtering in the range of 1 to 50 Hz. Subsequently, we extract samples from the raw signals, temporally aligning the data to a range spanning from -200 to 800 ms relative to the onset of the stimulus. Afterward, Following the baseline correction, where we subtract the mean of a baseline period (-200 to 0 ms) from the entire sample data points to mitigate drift effects like DC offsets [27]. Finally, we employ a peak-to-peak rejection method to eliminate large artifacts attributed to eye or muscle movements from contaminating the EEG data.
@@ 3.3 Feature Extraction
To compute AR coefficients, our system fits AR models to the pre-processed samples, employing a one-second duration of time series data [23]. The estimation of these AR coefficients utilized the Yule-Walker method [69]. Determining the ideal order for AR modeling poses a challenge, as higher orders escalate computational demands, while exceedingly low orders inadequately represent the signal [70]. To address this, our system allows users the flexibility to define their preferred AR order, enabling a tailored approach aligned with their specific requirements.
@@
Furthermore, NeuroIDBench implements a deep learning based approach using a TNN with triplet loss function to transform time-series EEG data into compact brain embeddings. Building on the work of Fallahi et al. [27], our TNN structure integrates three Convolutional Neural Network (CNN) branches, each featuring five convolutional layers, as illustrated in Figure 3. These branches received 1-second samples as input, structured in a two-dimensional array where rows correspond to channel indices, and columns represent discrete-time measurements. After each convolutional layer, we applied an average pooling layer to reduce the input vectors’ dimensionality while preserving the unique characteristics of each brainwave.
Figure 3: The CNN’s network architecture within the Twin neural Network is designed to produce a condensed 32-bit embedding from brain data samples, serving as an efficient method to compute latent representations of inputs [27].
@@ 3.4 Evaluation
While most EEG studies assess brainwaves in a single-session scenario [72, 27, 23, 29, 34, 35], it is crucial to examine EEG robustness over time. A realistic and often overlooked approach is the multi-session scenario. To highlight this gap, we provide both single-session and multi-session evaluation scenarios in our benchmark tools as follows
@@
For the known attacker scenario, we implemented a Stratified cross-validation approach with k=4 to divide the single-session data into training and testing sets. This choice ensures the representation of features from both classes in both training and testing data across each fold. We take the step of eliminating users with fewer than four samples from the datasets to guarantee sufficient samples for both training and testing [23]. Conversely, the unknown attacker approach adopts the GroupKFold cross-validation strategy with k=4, where grouping is based on SubjectID, resulting in non-overlapping training and testing sets for users in each cross-validation round. Evaluation metric results are aggregated across all folds, providing averaged and comprehensive reports. In datasets like COGBCI and Lee2019, which had multiple sessions, we evaluate each session separately, subsequently averaging the results from all the sessions as single-session evaluation results. Multi-session evaluation, described later in this section.
Additionally, we actively apply z-scaler normalization to both training and testing data to address potential issues arising from variations in feature scales, ensuring equitable contribution from all features during model training [73]. This approach aligns with established best practices in machine learning [74], reinforcing the reliability and generalization capability of our EEG-based authentication system.
Multi-Session Evaluation: The multi-session evaluation methodology was applied to two EEG datasets, namely COGBCI [49] and Lee2019 [61]. COGBCI consists of three sessions separated by day intervals, while Lee2019 includes two sessions separated by week intervals. A cross-validation strategy was employed to address overfitting. This involved dividing the attacker subjects for the shallow classifiers and dividing the subject of latent space learning and evaluation.
@@ 3.5 Analysis
Proper metrics are essential for comparing different methods from the viewpoint of uniqueness. We have chosen to report EER (Equal Error Rate) and FNMR (False Non-Match Rate) at various FMR (False Match Rate) levels, as recommended by the biometric authentication community [75, 76] and international standards [77, 78]. The FMR indicates the likelihood of an unauthorized user being mistakenly authenticated by the system, reflecting the system’s vulnerability to a zero-effort attack. The goal is to achieve a low FMR while maintaining an acceptable FNMR. A high FNMR can result in legitimate users being repeatedly denied access, negatively impacting the device’s usability. Therefore, it is crucial to balance FMR and FNMR to ensure robust security without compromising user experience. The EER represents the point, threshold, where the FMR and FNMR are equal.
@@ 4.1 Preprocessing
Figure 4 presents the EER results for the sample duration experiment. The results indicate that longer samples may enhance outcomes in shallow classifiers, but no significant improvement was observed in the case of Twin neural networks. It is important to note that for longer samples (e.g., 2 seconds), it was necessary to downsample the sample to mitigate the increased data dimensionality, a step taken due to hardware limitations as suggested in source code of BrainNet [27] paper as well.
@@ 4.3 Known and UnKnown Attacker Approach
The known attacker model represents an unrealistic scenario; however, given its existence in some previous studies, we investigated the performance gap between known and unknown attacker scenarios. Figure 7 displays the EER for these two scenarios, based on a sample duration of 1 second to ensure comparability with previous work. This analysis was conducted without sample rejection and utilized PSD combined with AR order 1 as the feature extraction method for shallow classifiers.
@@
For the practical application of brainwave authentication, it is essential to develop a model capable of handling multi-session authentication. However, there is a notable limitation in the availability of datasets for multi-session studies. Our investigation focuses on comparing single-session and multi-session authentication to understand the performance gap between these two approaches. We expected an increase in EER for multi-session authentication, as it represents a more complex challenge. This increase in difficulty is attributed to additional noise factors inherent in multi-session settings. Additionally, variations in EEG electrode placement across sessions and possible changes in brain states add to the complexity, making multi-session authentication a more challenging task compared to single-session scenarios.
@@
The lower performance of the deep learning approach compared to shallow classifiers in multi-session scenarios may result from the limited amount of data available for training feature extraction models. These models face complex issues across different sessions, such as changes in hair length, emotional states, and electrode positions, which complicate effective feature extraction with limited data. In contrast, classical feature extraction methods like PSD or AR require less data by relying on expert knowledge and do not need data for learning features. However, with more data, we expect twin neural network models to improve and potentially outperform shallow classifiers.
@@ 4.5 Shallow Classifiers vs Similarity-Based Approach
BrainInvaders15a Single-Session FNMR at 1% FMR 19.87 71.26 43.39 31.55 0.69 2.06 2.12
FNMR at 0.1% FMR 34.62 83.22 58.55 38.77 1.47 4.54 3.70
@@
ERPCORE_N400 Single-Session FNMR at 1% FMR 22.70 35.83 25.67 76.06 1.96 7.54 1.17
@@
ERPCORE_P300 Single-Session FNMR at 1% FMR 26.65 39.28 30.03 59.55 4.36 10.47 4.30
@@
Huebner_LLP Single-Session FNMR at 1% FMR 26.48 8.41 5.06 34.01 0.31 0.51 0.02
@@ 4.6 Comparison with Related Work
The comparison of our results will initially focus on the results published by Arias et al. [24] and Fallahi et al. [27]. This is due to the establishment of our benchmark, which is based on their implementations. Moreover, both studies utilized the ERPCORE [36] P300 and N400 datasets, similar to our approach. Arias et al. [24] investigate the performance of different shallow classifiers within single-session brainwave authentication. The results indicate 1.9% EER for N400 and 3% EER for P300 paradigm with a sample rejection of 120 µV and RF classifier, we observe the same results where we achieved 2.9% EER for P300 and 1.13% EER to N400 paradigm with the sample rejection of 100 µV and RF classifier. Also, Fallahi et al. [27] acquired 2.01% and 1.37% respectively for P300 and N400 ERPCORE and 0.14 for P300:bi2015a dataset [79], where we obtained %1.53, 0.83% , and 0.43 respectively. As can be observed the results are fairly similar and trends are the same. The small difference in results could be from a different random seed, version of libraries, and some evaluation parameters like epoch rejection rate for shallow classifiers and number of epochs for twin neural networks. To enhance the reproducibility of our result, we provided a docker container in the GitHub of NeuroIDBench.
@@
Preprocessing: Our results about the impact of sample duration corroborate prior research on shallow classifiers, which indicates that longer sample durations can result in reduced EER or increased accuracy [24, 80, 81, 82]. However, the current version of the twin neural network requires modification to effectively utilize longer samples.
@@
Feature Extraction: The results of our feature extraction analysis demonstrate that PSD features notably outperform AR features, corroborating findings from Huang et al. [83]. However, there is a lack of consensus in the literature regarding the optimal order for AR features. Studies by Arias et al. [23] and Brigham [84] suggest that lower AR orders yield better results, whereas research by Kaewwit et al. [85] and Zhang et al. [70] indicates improved outcomes with higher orders. Our findings align with the former, showing that lower-order AR features are more effective in 8 out of 9 datasets analyzed.
@@
Known and Unkown attacker: In exploring scenarios involving known and unknown attackers, Wu et al. [86] conducted research but found no notable performance decline in the unknown attacker scenario, and even reported an improvement. Similarly, Panzino et al. [87] studied the same issue and observed only a 5% decrease in accuracy. Contrarily, Arias et al.[23] and Fallahi et al.[27] reported an increase in Equal Error Rate (EER) ranging from 10% to 17.5 times. Our results indicate that the unknown attacker scenario is consistently more effective than the known attacker scenario, although the extent of this difference varies across different datasets and models. An average calculation across nine datasets and seven models shows that the EER for unknown attackers is 60.2% higher than known attacker scenarios.
@@
Single and Multi-session: Multi-session analysis is crucial for the practical application of brainwave authentication, yet it presents heightened challenges compared to single-session scenarios. In multi-session environments, the availability of public datasets is notably scarcer, and the practice of open-source sharing becomes more constrained. Results from various studies, such as Maiorana [30], demonstrate reasonable EER as low as 4.8%. Seha et al. [88] also reported a marginal difference of less than 0.5% EER between single and multi-session scenarios under specific feature extraction methods and LDA classifiers. Furthermore, Wu et al. [86] observed a 1.3% improvement in false rejection rate performance in the second session compared to the first session in single-session scenarios. In contrast, our research reveals a significant disparity between single-session and multi-session results, with our best outcome showing an 11.99% EER. This finding aligns with Huang et al. [83], who also reported a substantial difference between within-session and cross-session evaluations. However, it is important to note that the studies by Maiorana [30], Seha et al. [88], and Wu et al. [86] do not provide public datasets or open-source code for other researchers to utilize and build upon. Therefore, there is an urgent need for more open-source and publicly available datasets in multi-session studies in this field. In fact, understanding the exact state of this challenge is crucial to determine what steps are necessary to advance the field.
@@
Shallow Classifiers vs Similarity-Based Approach: The current trend in biometric authentication leans towards deep learning approaches. We observed the same trend in brainwave authentication where several papers based on TNN [27, 30, 34] or CNN [29] to improve authentication. However, a comparison between shallow classifiers and deep learning methods is essential. This is because the scarcity of data can hinder the training of deep learning networks, particularly in managing the high variation in brainwave data. Several studies [89, 27] demonstrate that deep learning approaches can outperform shallow classifiers. However, our large-scale benchmark emphasizes, despite the impressive results of TNN, they do not consistently surpass shallow classifies, Notably, in multi-session scenarios, TNNs underperform, potentially due to insufficient data to accommodate the high variability between sessions.
5 Limitations
The two main limitations observed while conducting research to develop our benchmark are the shortage of publicly available datasets and the evolving ethical and privacy concerns in the deployment of EEG-based biometric systems. These challenges underscore critical areas for future research and development to enhance the reliability and ethical assurance of EEG authentication.
Publicly Available Datasets: Currently available public datasets are not designed for authentication purposes. They often have a low number of subjects, usually fewer than 70, and mostly provide data from a single session, making it difficult to evaluate the robustness of EEG authentication over time (Table 2). Additionally, most datasets use medical-grade devices for data collection, complicating performance estimation in real-world scenarios. Therefore, there is a need to collect authentication datasets that include a larger number of subjects (over 500), with at least three sessions, and various environmental conditions, such as variations in noise, lighting, hair length, and emotional states. This will enable proper training and evaluation of models and allow for comparisons with well-known biometrics such as face recognition models.
Ethical and Privacy Implications: As EEG authentication research advances toward practical applications, increasing attention is being paid to the ethical aspects of collecting and using brain data for authentication. Previous studies have looked into privacy concerns from researchers [90, 91] and users [92, 93], identifying the need for protecting these data. The main issue is that EEG can reveal highly sensitive personal information, including emotional states [94], medical conditions [95], attention levels [96], and gender [97]. Therefore, collecting brain data raises the risk of harmful data breaches and introduces new possibilities for abuse or misuse. For example, an honest-but-curious authentication provider could perform unauthorized behavioral monitoring or exploitation of brain reactions for commercial purposes. This potential threat is aggravated by the lack of user awareness regarding the sensitivity of brain data when using commercial BCIs [98]. Beyond privacy-related concerns, there is a risk of bias in deployed EEG-based authentication systems if not trained with a diverse set of people. This type of issue has been observed in face recognition software that performs poorly when used by subjects who belong to underrepresented groups [99].
To address these concerns, we recommend the following general guidelines for ethical practice. In terms of data collection for research, researchers must adhere to the ethical standards established in the Menlo Report [100] and inform subjects about potential risks through consent forms. For real-world EEG authentication, it is crucial to develop robust template protection methods to prevent authentication providers from accessing raw EEG data, protecting against potential leaks and inferences. Systems should be trained to ensure fairness and reduce algorithmic bias, which entails the collection of inclusive datasets that capture the diversity of all potential users. Similarly, broad studies on the social acceptance and diverse user needs should be conducted to guide a responsible human-centered development of EEG biometrics. Finally, it is important to implement effective transparency mechanisms beyond unusable privacy policies to communicate to BCI users what the collected data and privacy risks are.